OpenStudy (unklerhaukus):

\[m\frac{\text dv}{\text dt}=mg-kv;\qquad v(0)=0\]

OpenStudy (unklerhaukus):

\[\frac{\text dv}{\text dt}+\frac kmv=g\]\[\qquad\qquad\qquad \text{let}\frac km=\omega_0^2\]\[\frac{\text dv}{\text dt}+\omega_0^2v=g\]

OpenStudy (unklerhaukus):

\[\mu(t)=e^{\int \omega_0^2\text dv}=e^{\omega_0^2 v}\] \[\frac{\text d }{\text dt}\left(v\cdot e^{\omega_0^2t}\right)=g\cdot e^{\omega_0^2 t}\]

OpenStudy (unklerhaukus):

\[v\cdot e^{\omega_0^2 t}=\int g\cdot e^{\omega_0^2 t}\text dt\]\[v\cdot e^{\omega_0^2 t}=ge^{\omega_0^2 }\int e^{t}\text dt\]\[v\cdot e^{\omega_0^2 t}=ge^{\omega_0^2}\left(t{e^t}+c\right)\]\[v(t)=g\left(t+ce^{-t}\right)\]\[v(0)=gc=0\]\[c=0\]\[v(t)=gt\]

OpenStudy (unklerhaukus):

is this correct?

OpenStudy (anonymous):

between lines 1 and 2 on reply 3, I don't think you can factor out omega like you did

OpenStudy (anonymous):

\[e^n * e^m = e^{n+m}\]

OpenStudy (anonymous):

\[\int\limits e^{t}\text dt\] = ??

OpenStudy (unklerhaukus):

e^t

OpenStudy (unklerhaukus):

i knew i was doing something wrong in the middle section

OpenStudy (anonymous):

Then how did t come with this in third step ??

OpenStudy (anonymous):

In fourth step sorry.. \[v\cdot e^{\omega_0^2 t}=ge^{\omega_0^2 }\int e^{t}\text dt\] After this ??

OpenStudy (anonymous):

I get: \[\frac{gm}{k} + C\]

OpenStudy (unklerhaukus):

\[\frac{\text dv}{\text dt}+\omega_0^2v=g\]\[\mu(t)=e^{\int \omega_0^2\text dt}=e^{\omega_0^2 t}\]\[\frac{\text d }{\text dt}\left(v\cdot e^{\omega_0^2t}\right)=g\cdot e^{\omega_0^2 t}\]

OpenStudy (anonymous):

Can't find anything wrong with that

OpenStudy (anonymous):

Then again it is 3:30 AM where I live

OpenStudy (unklerhaukus):

\[v\cdot e^{\omega_0^2 t}=\int\limits_0^t g\cdot e^{\omega_0^2 t'}\text dt'\]\[v\cdot e^{\omega_0^2 t}=ge^{\omega_0^2 }\int\limits_0^t e^{t'}\text dt'\]\[v\cdot e^{\omega_0^2 t}=ge^{\omega_0^2}\left(e^t+1\right)\]

OpenStudy (unklerhaukus):

is this better ? \[v(t)=g\left(1+ce^{-t}\right)\]\[v(0)=g(1+c)=0\]\[c=0\]\[v(t)=gt\]

OpenStudy (anonymous):

I don't think you can factor a: \[e^\omega \] out of the integral

OpenStudy (anonymous):

You can take out uncle rocks as that is the constant only..

OpenStudy (unklerhaukus):

\(m\) and \(k\) are constants

OpenStudy (anonymous):

Yes so omega will also be constant..

OpenStudy (anonymous):

but: \[e^{\omega t} \neq e^ \omega * e^t\]

OpenStudy (anonymous):

and in fact if you evaluate the two integrals you will see that they are different

OpenStudy (anonymous):

by a factor of omega

OpenStudy (anonymous):

or whatever we are using, omega^2

OpenStudy (anonymous):

Did not you get c= -1 instead of 0 @UnkleRhaukus

OpenStudy (unklerhaukus):

hmmmmmmm

OpenStudy (unklerhaukus):

im a little confused

OpenStudy (anonymous):

\[\large v\cdot e^{\omega_0^2 t}=\int\limits\limits_0^t g\cdot e^{\omega_0^2 t'}\text dt'\] \[\large v\cdot e^{\omega_0^2 t} = g \left| \frac{e^{(\omega_0)^2t'}}{(\omega_0)^2} \right|_{0}^{t}\]

OpenStudy (unklerhaukus):

\[v\cdot e^{\omega_0^2 t}=\int\limits_0^t g\cdot e^{\omega_0^2 t'}\text dt'\]\[v\cdot e^{\omega_0^2 t}=g\int\limits_0^t e^{\omega_0^2 t'}\text dt'\]\[v\cdot e^{\omega_0^2 t}=g\left.\frac {e^{\omega_0^2t'}}{\omega_0^2}\right|_0^t \]\[v\cdot e^{\omega_0^2 t}=g\frac {e^{\omega_0^2t}-1}{\omega_0^2} \]\[v(t)=g\left(1+e^{-t}\right)\]

OpenStudy (anonymous):

Sorry you cannot separate omega squared and t because they are in multiplication and not addition..

OpenStudy (anonymous):

\[\large v(t)=\frac{g}{(\omega_0)^2}\left(1-e^{-(\omega_0)^2t}\right)\]

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[v\cdot e^{\omega_0^2 t}=g\frac {e^{\omega_0^2t}-1}{\omega_0^2} \]\[v(t)=\frac{g}{\omega_0^2}\left(1-e^{-\omega_0^2t}\right)\]\[v(0)=\frac{g}{\omega_0^2}\left(1-1\right)=0\]\[v(t)={g}\frac{\left(1-e^{-\omega_0^2t}\right)}{\omega_0^2}\]

OpenStudy (anonymous):

Yes now put the conditions..

OpenStudy (unklerhaukus):

i though i did that?

mathslover (mathslover):

it seemed like a "tutorial" :)

OpenStudy (anonymous):

well distance traveled is \[ve^{\int\limits_{}^{}kt/m}=\int\limits_{}^{}{g}e^{\int\limits_{}^{}kt/m}dt+c\]

OpenStudy (unklerhaukus):

have i finished this question or what//?

OpenStudy (anonymous):

so\[{g{e}^{kt/m}\over{{k/m}}}+c = {{gm}\over{k}}{e}^{kt/m} +c\] so v \[v={{gm}\over{k}}+ce^{-kt/m} \] intial cond. \[0=v={{gm}\over{k}}+ce^{-kt/m} \]

OpenStudy (anonymous):

sorry my computer bugs out whenever i use the equation thing

mathslover (mathslover):

try to type that in notepad and then copy paste it here @noah_ochoa :)

OpenStudy (anonymous):

if x = distance traveled in t and \[c=-gm/t\]

OpenStudy (unklerhaukus):

\[v(t)=\frac {mg}k\left(1-e^{-\frac {kt}m}\right)\]

OpenStudy (anonymous):